Calculate the ph of the solution after the addition of the following amounts of 0.0576 M HNO3 to a 50.0 mL solution of 0.0750 Maziridine. The pKa of aziridinium is 8.04. d) 61.9 mL of HN a) 0.00 mL of HNO3 Number Number pH 6.70 pH 10.458 b) 9.00 mL of HNO e) Volume of HNO3 equal to the equivalence point Number Number pH 8.90 pH c) Volume of HNO3 equal to half 70.1 mL of HNO3 the equivalence point volume Number Number pH 8.40 pH IncorrectI really need help with c-f please
Answer #1

c)

in the half equivalence point

pKb =14- 8.04 = 5.96

then

pOH = pKb + log(BH+/B)

HB+ = B in half equivalence point so;

pOH = 5.96

pH = 14-5.96= 8.04

e)

HNO3 for equivlanece point

mmol of acid = mmol of base

Macid*Vacid = Mbase*Vbase

0.0576 * Vacid= 0.075*50

Vacid= 0.075*50/0.0576 = 65.104 mL of acid required

f)

this is mostly acidic, since there is excess H+ strong base

mmol of HNO3 = MV = 70.1*0.0576 = 4.03776 mmol of HNO3

mmol of base = MV = 0.075*50 = 3.75

mmol of hno3 excess = 4.03776 -3.75 = 0.28776 mmol of H+

V total =) 70.1 + 50 = 120.1 mL

pH = -log(h+) = -log(0.28776 /120.1) = 2.6205